what is the area of a rectangle with a diagonal length of 26m and height of 10m, round answer to nearest tenth
Answer by richwmiller(17219)

no need to round
if the diagonal is 26 and the height is 10 then we have a multiple of the 5,12,13 right triangle 10,24,26
so the base is 12*2=24
other wise 10^2+a^2=26^2
solve for a
a=24
10*24=240 sq m.